OpenStudy (anonymous):

Kevin picked 13 more apples than Andy. If Kevin picked 30 apples, how many did Andy pick?

OpenStudy (mathstudent55):

Kevin picked k apples Kevin picked 30 apples, so k = 30 Andy picked "a" apples Kevin picked 13 more apples than Andy k = 13 + a k = 13 + a 30 = 13 + a a + 13 = 30 Can you solve this equation?